### v2

# LLMs Are Biased Towards Output Formats! Systematically Evaluating and Mitigating Output Format Bias of LLMs

Abstract
We present the first systematic evaluation examining format bias in performance of large language models (LLMs). Our approach distinguishes between two categories of an evaluation metric under format constraints to reliably and accurately assess performance: one measures performance when format constraints are adhered to, while the other evaluates performance regardless of constraint adherence. We then define a metric for measuring the format bias of LLMs and establish effective strategies to reduce it. Subsequently, we present our empirical format bias evaluation spanning four commonly used categories—multiple-choice question-answer, wrapping, list, and mapping—covering $15$ widely-used formats. Our evaluation on eight generation tasks uncovers significant format bias across state-of-the-art LLMs. We further discover that improving the format-instruction following capabilities of LLMs across formats potentially reduces format bias. Based on our evaluation findings, we study prompting and fine-tuning with synthesized format data techniques to mitigate format bias. Our methods successfully reduce the variance in ChatGPT’s performance among wrapping formats from $235.33$ to $0.71$ (%$\%^{2}$).
1 Introduction
Figure 1: A MMLU example Hendrycks et al. (2021) with ChatGPT across different formats. In Case (1), the model can answer the question but fails to bold only the answer, hindering automatic evaluation. In Case (2), the model follows the format but produces an incorrect result. In Case (3), the model yields the correct answer and format. These show bias in ChatGPT’s performance across formats.
To unlock the full potential of automating real-world applications, state-of-the-art large language models (LLMs) (Brown et al., 2020; Chowdhery et al., 2022; OpenAI, 2022; Touvron et al., 2023) are increasingly leveraged to tailor outputs to specific task formats. This powerful approach has driven advancements across domains including medicine Thirunavukarasu et al. (2023); Clusmann et al. (2023), data analysis Cheng et al. (2023); Liu et al. (2023), and even evaluating models themselves Chiang and Lee (2023); Chang et al. (2024). Employing LLMs in such applications heavily depends on not only their format-following capability but also high-quality results within formats.
While many studies, including those listed above, have utilized LLMs to output in specific formats, understanding their format capabilities is critical yet has received limited attention. Recently, Zhou et al. (2023) and Xia et al. (2024) introduced benchmarks assessing LLM format-following proficiency. However, these studies neglect deeper insights into how these formats impact model performance, which is the ultimate concern for industrial and practical applications. Given numerous formats recently introduced across tasks and models, assessing this aspect is essential for business yet challenging. Evaluation can be ambiguous and often overlook cases where models provide correct answers but are formatted wrong (Case (1) in Fig. 1).
Bridging these gaps, we conduct the first systematic evaluation of the format bias of LLMs. Our study attempts to answer the research questions:
How can we systematically and accurately assess format bias in the performance of LLMs, and to what extent are they biased?
To fairly assess bias in model performance across formats, it is crucial to evaluate all scenarios depicted in Fig. 1. Nonetheless, Case (1) is challenging to automatically measure, requiring costly human investigation. Therefore, we propose a reliable estimator for evaluating LLM performance under format constraints without human intervention by considering format-following scores. We start by redefining LLM evaluation metrics into two distinct classes to construct the estimator, as detailed in Section 3.1. Accordingly, we define a metric to quantify format bias in LLMs and establish criteria for evaluating methods that successfully mitigate this bias (Section 3.2). Based on these formulations, we present our format evaluation framework, comprising of the widely-utilized categories of multiple-choice question–answer (MCQ; Section 5.1), wrapping (Section 5.2), list (Section 5.3) and mapping formats (Section 5.4).
Across $15$ widely-used formats, our evaluation with zero-shot and zero-shot chain-of-thought prompting Kojima et al. (2022) on eight question-answering and reasoning tasks reveals substantial performance and format-instruction following inequalities. To address this, we examine prompting and fine-tuning using synthesized format data techniques that work for both open- and closed-source LLMs. Our study validates that enhancing LLMs’ capabilities to follow format instructions potentially mitigates format bias: (1) Prompting with demonstrations and (2) Repeating format instructions substantially alleviates this bias. Moreover, we investigate (3) Synthesizing limited format data based on our evaluation results for fine-tuning. Our approaches significantly decrease ChatGPT performance variance across wrapping formats from $235.33$ to $0.71$ (%$\%^{2}$) on MMLU Hendrycks et al. (2021). Our key contributions are:
  1. 1.
    We introduce the first systematic framework [^1] Our codes and data will be made publicly available at link. to evaluate format performance bias in LLMs.
  2. 2.
    A large-scale evaluation spanning $15$ formats, $8$ tasks, and $4$ models revealing substantial LLM performance variance across formats.
  3. 3.
    The development of $3$ novel prompting and fine-tuning methods to mitigate this bias.
2 Related Works
Large language models (LLMs) have shown remarkable proficiency in formatting outputs to meet human expectations. Such formats include markdown for lists and pointers (Achiam et al., 2023), code blocks (Gur et al., 2023), and integrate tags, or LaTeX for scientific texts (Singh et al., 2023; Wang et al., 2024). Given the rising importance of formatting capabilities in LLMs, recently, format-following benchmarks have been developed for assessing LLMs’ adherence to specified formats (Zhou et al., 2023; Xia et al., 2024; Chen et al., 2024; Macedo et al., 2024; Liu et al., 2024). However, these studies only evaluate format-instruction following capabilities. Our research further assesses LLM performance across different formats, uncovering significant format bias in various tasks and models. We also acknowledge the concurrent work by Tam et al. (2024), which examines the impact of format restrictions on LLM performance. However, unlike our approach, they do not disentangle evaluation metrics under format constraints and only evaluate $3$ structured formats, substantially fewer than our study.
3 Output Format Evaluation Framework
3.1 Theoretical Analysis: Format Evaluation
Automatic evaluation of LLMs in question-answering and reasoning tasks mainly relies on rule-based extraction to identify final answers from generated texts Guo et al. (2023). Within format constraints, determining the model’s true performance, which is our focus, can be ambiguous and inaccurate, as correct responses might be overlooked due to format discrepancies (e.g., Case (1) in Fig. 1). To address this, we propose redefining these rule-based evaluation metrics to reliably, transparently and accurately measuring the LLM performance given formats restrictions.
Notations.
Suppose that we are interested in evaluating an LLM $\mathcal{M}$ on a task $T$ using an evaluation metric $E$ (such as “Accuracy”) under a format constraints $C$ (such as “Bold the final answer.”) on $n$ samples with the ground-truth answers $\{y_{1},...,y_{n}\}$ and raw generated answers $\{\hat{y}_{1},...,\hat{y}_{n}\}$, where $y_{i},\hat{y}_{i}\in\mathcal{Y}$ $\forall i$ with $\mathcal{Y}$ being the answer token sequence space. We denote $F_{C}$ as the binary format-following evaluation function of $C$:
$F_{C}(\hat{y}_{i})=\begin{cases}1,&\text{if}\ \hat{y}_{i}\ \text{satisfies}\ C.\\ 0,&\text{otherwise.}\end{cases}$ (1)
From Eq. 1, we define the Format Instruction-following (FI) Score, denoted as $FI_{C}$, as the percentage of generated outputs satsisfying $C$:
$FI_{C}=\frac{\sum_{i=1}^{n}F_{C}(\hat{y}_{i})}{n}\cdot 100$ (2)
Prior studies extensively focus on evaluating $FI_{C}$ Zhou et al. (2023); Xia et al. (2024). Our work further targets evaluating the performance of LLMs given the format constraints $C$. Under $C$, we denote $Ext_{C}()$ as the rule-based answer extractor (or a mixture of extractors) to extract the final answer from $\hat{y_{i}}$ for comparing it with $y_{i}$. We define: two evaluation scores based on $E$:
Definition 3.1 (Systematic Evaluation Score ($SysE$)).
$SysE=\frac{1}{n}\sum_{i=1}^{n}(E(y_{i},Ext_{C}(\hat{y_{i}})).F_{C}(\hat{y_{i}}))$ (3)
Essentially, $SysE$ quantifies the performance of $\mathcal{M}$ on task $T$ based on the generated answers that meet the format constraints $C$. For example, in Fig. 1, Case (1) yields a $SysE$ score of 00, while Case (3) achieves $1$. This also shows that $SysE$ may not accurately reflect the actual performance of $\mathcal{M}$ on $T$, because $Ext_{C}()$ may fail to extract the final answers from (correct) answers dissatisfying $C$ (e.g., Case (1) in Fig. 1). We define the True Evaluation Score to address this. Assume that we have an oracle extractor function $OracExt_{C}()$ that can extract the final answer from $\hat{y_{i}}$, regardless of whether $\hat{y_{i}}$ fulfills $C$, we have:
Definition 3.2 (True Evaluation Score ($TrueE$)).
$TrueE=\frac{1}{n}\sum_{i=1}^{n}E(y_{i},OracExt_{C}(\hat{y_{i}}))$ (4)
$TrueE$ measures the performance of $\mathcal{M}$ on task $T$ across all generated answers given the format constraints $C$, regardless of format satisfaction. In Fig. 1, both Cases (1) and (3) achieve a true accuracy of $1$. This score is crucial for assessing the true performance of LLMs given the format.
Prior studies do not clearly differentiate between $SysE$ and $TrueE$. In practice, measuring $TrueE$ is challenging because $OracExt_{C}()$ is unavailable. While researchers typically employ a mixture of methods to extract answers, this approach encounters two severe issues. First, these mixture-of-method extractors can be complex, unreliable, and often impractical for large-scale experiments with diverse formats like ours. Second, designing them to be reliable for complex formats such as medical reports can be impossible due to the countless potential errors. Another alternative is to assign a default value to $Ext_{C}(\hat{y_{i}})$. While this can temporarily avoid cases $\mathcal{M}$ fails to fulfill $C$, this is an incorrect practice since the default value may not be the actual output. Reliably measuring $TrueE$ often requires human investigation (Lin et al., 2022) or the fine-tuning of evaluation models as scorers (Yang et al., 2024), both of which are costly.
Nevertheless, $TrueE$ is crucial for a fair evaluation of LLM performance bias across formats. Therefore, we propose a simple estimator of $TrueE$, denoted as $EstTrueE$:
$EstTrueE=\begin{cases}SysE.\frac{100}{FI_{C}},&\text{if}\ FI_{C}\neq 0.\\ 0,&\text{otherwise.}\end{cases}$ (5)
When $FI_{C}=0$, estimating $EstTrueE$ becomes impossible. $EstTrueE$ enables the fair format bias evaluation because normalizing $SysE$ by $FI_{C}$ prevents skewing comparisons of how different formats affect the LLM due to $FI_{C}$. It is especially useful for large-scale experiments since it is fully automatic. Let the $EstTrueE$ margin of error be $\epsilon$ with a confidence interval $1-\alpha$ and $S_{C}=n\cdot FI_{C}$ as #generated answers satisfying $C$.
Theorem 3.1 (Reliability of $EstTrueE$).
$EstTrueE$ is consistent. Moreover, $EstTrueE$ is reliable if and only if:
$FI_{C}\geq\frac{1}{1+n\cdot\left(\frac{\epsilon}{v\cdot s}\right)^{2}}$ (6)
Moreover, we have:
$\lim_{{FI_{C}}\to 100}{EstTrueE}={TrueE}$ (7)
where $s^{2}$ is the sample variance of evaluation scores of generated answers satisfying $C$ and $v=t_{\alpha/2,S_{C}-1}$ is the critical value from the t-distribution with $S_{C}-1$ degrees of freedom.
In summary, we have proposed a consistent estimator $EstTrueE$ of the true performance of LLMs measured by metric $E$ under the format constraints (Def. 3.2). This estimator is essential because it: (1) ensures transparent and fair LLM performance evaluation across different formats; and (2) supports large-scale format bias evaluation. Note that a high score $EstTrueE$ is only reliable iff $FI_{C}$ is high enough (Theorem 3.1). Henceforth, unless otherwise specified, $EstTrueE$ is our primary metric for measuring model performance given format constraints. The proof of Theorem 3.1 is in Section B.1.
3.2 Theoretical Analysis: Format Bias
This section defines the metric to quantify format bias and outlines the criteria to mitigate such bias.
Bias measurement.
To measure the format bias of the LLM $\mathcal{M}$ across $k$ formats $F_{o}=\{C_{1},\ldots,C_{k}\}$, we define a single metric, $BiasF_{o}$, as the variance of $EstTrueE$ scores over these $k$ formats, denoted as $\{EstTrueE_{1},\ldots,EstTrueE_{k}\}$. Let $\mu_{EstTrueE}=\frac{1}{k}\sum_{i=1}^{k}EstTrueE_{i}$ represent the mean $EstTrueE$ score. Then:
$BiasF_{o}=\frac{1}{k}\sum_{i=1}^{k}(EstTrueE_{i}-\mu_{EstTrueE})^{2}$ (8)
Realiability of $BiasF_{o}$.
By Eq. 8, the lower $BiasF_{o}$ is, the less format-$F_{o}$-biased $\mathcal{M}$ is, suggesting a criterion for mitigating output format bias. However, $BiasF_{o}$ is an estimator based on the estimators $EstTrueE_{i}$. Therefore, to enhance the reliability of $BiasF_{o}$, it is also necessary to improve the reliability of $EstTrueE_{i}$ by increasing $FI_{C_{i}}$ $\forall i$ (Theorem 3.1). Therefore, we propose two necessary criteria for an effective method to mitigate format bias in LLMs: (i) Minimize bias metric: reducing $BiasF_{o}$, indicating less format-$F_{o}$-bias in $\mathcal{M}$; (ii) Increase the format-following scores for all formats: ensuring the reliability of $BiasF_{o}$ by increasing the $FI$ scores across all the formats: $\{FI_{C_{1}},...,FI_{C_{k}}\}$ (Eq. 2).
3.3 Formats for Evaluation
We establish $4$ format categories for evaluation consisting of $15$ formats introduced by prior practices:
(i) Multiple-choice question (MCQ) answer (Section 5.1).
where LLMs answer questions by selecting from provided choices, presented as either a (1) Character identifier Robinson and Wingate (2023); or (2) Choice value Chen et al. (2023).
(ii) Wrapping (Section 5.2).
where LLMs must enclose the final answer within the two characters, which is crucial for automatic evaluation to isolate the final answer from reasoning thoughts. We focus on evaluating $7$ widely used wrapping strategies: (1) Special character Gur et al. (2023); (2) Bolding (Zhou et al., 2023); (3) Italicizing (Zhou et al., 2023); (4) Double brackets (Luo et al., 2024); (5) Double parentheses; (6) Placeholder (Wang et al., 2024); (7) Quoting (Zhou et al., 2023).
(iii) List (Section 5.3).
where the output of LLMs is a list of elements. We investigate $4$ formats representing lists: (1) Python list Do et al. (2025); (2) Bullet-point list Liu et al. (2024); (3) List of elements separated by a special character “[SEP]” Boucher (2023); and (4) List of elements arranged on separate lines Mishra (2023).
(iv) Mapping (Section 5.4).
where LLMs are employed to output dictionaries or maps. We focus on two ubiquitously used mapping structures: (1) Python dictionary/JSON (JavaScript Object Notation) (Baumann et al., 2024) and (2) YAML (Yet Another Markup Language) (Goel et al., 2023).
Format-instruction following.
We introduce Appx.-Alg. 1, a rule-based heuristic to determine the format-instruction following function $F_{C}$ (Eq. 1) for our benchmarked formats. It calculates the binary FI score by verifying that the generated output includes the specified formatting tokens and that the extracted final answer matches the expected type. It is highly extendable to other formats (Appendix A).
4 General Experimental Setups
Benchmarks.
For MCQ bias evaluation (Section 5.1), we select two datasets: MMLU (Hendrycks et al., 2021) and BBH (Suzgun et al., 2023). For MMLU, we randomly choose 27 subcategories. For BBH, we select the sports_understanding category following Gupta et al. (2024). For wrapping bias assessment (Section 5.2), in addition to MCQ benchmarks, the following datasets are experimented: GSM8K (Cobbe et al., 2021) for reasoning, FairytaleQA (Xu et al., 2022) for narrative comprehension, and HotpotQA (Yang et al., 2018) for multi-hop reasoning. For list bias investigation (Section 5.3), we use SciDocsRR (Muennighoff et al., 2023), a scientific document ranking task as the order list generation task, and SemEval 2017 (Augenstein et al., 2017), the keyphrase extraction task as the unordered list generation. For mapping bias examination (Section 5.4), we utilize a document-level information extraction task named SciREX (Jain et al., 2020) by synthesizing three extraction difficulty levels: easy (extracting from $1$ sentence for $1$ category), medium ($3$ sentences, $2$ categories), and hard ($5$ sentences, $4$ categories). For all benchmarks except MCQ, we sample $200$ points for evaluation Bai et al. (2024).
Models.
We select both open- and closed-source LLMs for our evaluation: Gemma-7B-it (Team et al., 2024), Mistral-7B-it-v0.2 (Jiang et al., 2023), and Llama-3.1-8B-it (Dubey et al., 2024) for open-source as they are among state-of-the-art open-source LLMs; ChatGPT (gpt-3.5-turbo-0125) for closed-source as this premier chatbot possesses superior instruction-following ability. Our purpose is not to reproduce the models’ performance, but to show the bias.
Metrics.
Following our discussion in Section 3.1, we disentangle Accuracy (Acc) for MMLU and BBH (Guo et al., 2023); F1 for GSM8K, HotpotQA, FairytaleQA; and Mean Average Precision (MAP) for SciDocsRR (Muennighoff et al., 2023) and we report the metrics $EstTrueAcc$, $EstTrueF1$, $EstTrueMAP$ (Eq. 5) in the main text. For metrics’ reliability, we set $\alpha$ = $\epsilon=5\%$.
Prompting baselines.
Our focus is on two widely used prompting baselines: (1) Zero-shot (ZS) prompting and (2) Zero-shot Chain-of-Thought (ZS-CoT) prompting (Kojima et al., 2022). For the ZS baseline, we instruct LLMs to answer the question with the prompt “Answer the following question…” followed by the suffix “without any explanation”. For ZS-CoT, we use the suffix “step-by-step” instead. For the ZS-CoT experiments in Sections 5.1, 5.3 and 5.4, LLMs are instructed to wrap the final answer by “<ANSWER>” and “</ANSWER>” tokens to distinctly isolate it from the reasoning chains (see Tab. 1 for the wrapping instruction). We use this wrapping method since our experiment in Section 5.2 shows that it achieves the highest instruction-following score on average across LLMs. Detailed prompts are provided in Appendix E. We average the performance under two prompting methods to report in the main text.
5 Format Evaluation Experiments
Overall, we find that: (1) Models show substantial format-following bias across formats for all benchmarks; (2) For all models and datasets, significant performance bias exists across formats; (3) $78.30\%$ of the $EstTrue$ results are reliable (with $70\%$ for MCQ, $82.5\%$ for wrapping, $67.19\%$ for list, and $77.08\%$ for mapping formats) highlighting significant weaknesses of LLMs in following format instructions. We dive into (2) for every format as it is our main focus, (1, 3) are discussed in detail in Sections C.1, C.2, C.3 and C.4.
5.1 Experiments on MCQ Format
Setups.
We investigate the bias of LLMs towards different MCQ output formats. We assess two formats as introduced in Section 3.3: (1) Character identifier and (2) Choice value. For example, if the choice is “[A. Yes, B. No]”, then the character identifier can be “A/B”, while the choice value can be “Yes/No”. We exclude the format combining the character identifier and choice value (such as “A. Yes”) from our evaluation because instructing LLMs to output this format can be non-trivial and require manual effort to craft instructions tailored for different models. To ensure that LLMs understand the “Character identifier” and “Choice value” as we expect, we add a contrastive format requirement to the prompts (e.g., “without any textual description” for the “Character identifier” prompts).
Results.
Fig. 2 provides a synopsis of our evaluation results, with numerical values shown in Appx.-Tab. 2. From Fig. 2-left, we observe that Mistral possesses the highest disparity between the two MCQ answer formats, with $58.69\%$ accuracy on average for character and only $4.22\%$ for textual value. Additionally, despite ChatGPT often being regarded as one of the most robust LLMs, it shows a significant performance difference between the two formats ($19.03\%$). Overall, LLMs are heavily biased towards outputting character identifiers. Requiring them to generate the choice’s value causes notable performance drops on most models.
Figure 2: Average estimated true accuracy (Section 3.1) results of MCQ benchmarks across models (left) and datasets (right) showing performance bias of LLMs across formats.
From Fig. 2-right, we notice that the models exhibit higher bias on BBH, which appears to be an easier benchmark than MMLU. We attribute this to the small size of BBH, which makes the performance more sensitive to format variations.
Why such bias?
We hypothesize the root cause of the significant performance bias across different formats is the format token bias of LLMs. The non-uniform distribution of FI scores among formats suggests that the models assign probabilities to format instructions differently based on their training data. This leads to varying prior assignments of probabilities to specific tokens, causing final predictions non-uniformly distributed across formats. This hypothesis is supported by our simple fine-tuning with formatted data, which familiarizes LLMs with format instructions relatively equally leading to a drastic format bias reduction (Section 6). This emphasizes the necessity of more research in fine-tuning LLMs to reduce format bias and raises concerns about the reliability and reproducibility of recent studies using varied formats.
5.2 Experiments on Wrapping Format
Wrapping type (start, end) Prompt: Wrap your final answer…
Special char. (<ANSWER>, </ANSWER>) by <ANSWER> and </ANSWER>.
Bolding (**, **) in bold by enclosing it with double asterisks.
Italicizing (*, *) in italics by enclosing it with single asterisks.
Brackets ([[, ]]) using double square brackets.
Parentheses (((, ))) using double parentheses.
Placeholder None by filling in the placeholder below:
“So the answer is: [placeholder]”
Quoting (”””, ”””) using triple double-quotation marks.
Table 1: Wrapping “start” and “end” tokens with instructions.
Figure 3: Average estimated true Accuracy (MCQ) and F1 (GSM8K, HotpotQA, FairytaleQA) scores (Section 3.1) across models (left) and across benchmarks (right), showing performance bias of LLMs across $7$ widely used wrapping methods.
Setups.
We study LLM bias towards $7$ wrapping methods: (1) Special character; (2) Bolding; (3) Italicizing; (4) Brackets; (5) Parentheses; (6) Placeholder; (7) Quoting, detailed in Tab. 1. We evaluate LLM performance across formats on the MMLU, BBH, GSM8K, FairytaleQA, and HotpotQA.
Results.
Fig. 3 outlines an overview of our evaluation outcomes with results in Appx.-Tab. 6. From Fig. 3-left, we see that Llama exhibits the highest bias towards different formats (with a $BiasF_{o}$ value of $74.86\%^{2}$; see Appx.-Tab. 7), while ChatGPT performs the best. Notably, for “Quoting” and “Parenthesis”, the Gemma follows instructions only about $0-4\%$ yielding nearly zero performance, highlighting its critical weaknesses. Among the 7 formats, “Placeholder” ($37.15\%$) proves to be the most effective wrapping output format, while “Quoting” ($24.58\%$), “Parenthesis” ($28.57\%$) are among those that achieve the lowest performance.
From Fig. 3-right, models exhibit bias across all tasks, with the lowest on MMLU ($16.58\%^{2}$; see Appx.-Tab. 7) possibly because the models already performed relatively well on it, and the highest on BBH ($54.58\%^{2}$), the challenging task without train data. This demonstrates the pervasive presence of wrapping bias in LLMs.
Why such bias?
The format token bias of LLMs as explained in Section 5.1 is also our hypothesis. Specifically, we found the low performance of the “Quoting” and “Parenthesis” because, in generation tasks, models often wrap (via quoting/parenthesizing) not only the final answer, as instructed, but also parts of the context (e.g., “‘The answer is 3.”’), leading to poor F1 scores. Moreover, Gemma completely ignores the above format instructions, resulting in 0% FI scores, which also contribute to the low average estimated F1 scores. These strongly indicate the presence of format token bias.
5.3 Experiments on List Format
Setups.
We explore the bias of LLMs in generating lists following $4$ formats: (1) Python list, (2) Bullet-point list, (3) Character-separated list, and (4) Newline-separated list. We evaluate the models on two list generation tasks: (i) Unordered list, using the keyphrase extraction task on the SemEval 2017 dataset, and (ii) Ordered list, using the document ranking problem on the SciDocsRR task.
Results.
Fig. 4 displays the key findings of our evaluation across models and datasets with numerical results in Appx.-Tab. 10. From Fig. 4-left, we notice that Mistral exhibits the most bias, with the $BiasF_{o}$ value of $353.80\%^{2}$. In contrast, ChatGPT and Gemma show much lower bias, with values of $7.08\%^{2}$ and $1.32\%^{2}$, respectively. Of the four formats, the “Python” and “Newline-separated” formats yield the highest performance, likely due to models trained extensively on code data. Conversely, the “Bullet-point list” format results in the lowest performance, particularly for Mistral, highlighting the inherent bias for such formats.
Figure 4: Average $EstTrueF1$ (SemEval2017) and $EstTrueMAP$ (SciDocsRR) (Section 3.1) across models (left) and benchmarks (right) showing performance difference of LLMs across $4$ widely used list formats.
The performance bias is regardless of the task as plotted in Fig. 4-right, with the highest $BiasF_{o}$ value of $54.12\%^{2}$ on the order list generation task SciDocsRR, and significantly lower ($31.86\%^{2}$) on SemEval2017 task. The high bias in the SciDocsRR task is because Mistral and Gemma mostly failed to perform this task following the “Bullet” and “Special character” list formats while excelling in solving it following the other formats.
Why such bias?
We attribute the bias to the format token bias (Section 5.1). Since the models were extensively trained on code data, they excel in solving code-related instructions. In contrast, “Bullet-point” and “Special character” lists are much less common. One interesting case is Gemma where it performed worse on generating “Python” lists compared to “Bullet-point” lists. Our analysis suggests that Gemma misinterprets the format instruction as a coding request, generating Python code programs instead of an answer in a Python list, suggesting Gemma was predominantly trained on code data.
5.4 Experiments on Mapping Format
Setups.
We examine the performance bias of LLMs on two mapping formats as discussed in Section 3: (1) Python dictionary/JSON; (2) YAML. We preprocess the SciREX task Jain et al. (2020) as described in Section 4 into three extraction levels: (i) Easy (1 sentence, “Task” category); (2) Medium (3 sentences, “Task, Method”); (3) Hard (5 sentences, “Task, Method, Material, Metric” categories).
Results.
Fig. 5 illustrates a summary of our evaluation with numerical details in Appdx.-Tab. 14. From Fig. 5-left, Gemma is the most biased, with a performance gap of $16.51\%$ between the two formats, followed by Mistral with a $16.07\%$ gap. ChatGPT and Llama, however, are relatively robust against format variations. On average, JSON performs significantly better than YAML for mapping, likely because more JSON data is used to train models due to its popularity.
Figure 5: Average estimated true F1 scores (Section 3.1) across models (left) and benchmarks (right) showing performance bias of LLMs across $2$ widely used mapping formats.
From Fig. 5-right, extracting $4$ categories in the Hard task shows the largest performance gap between mapping formats. Surprisingly, the Medium task displays the least bias, likely because models perform best in this task.
Why such bias?
The bias is attributed to the format token bias (Section 5.1). While Mistral excels in generating JSON, it and Gemma struggle with YAML. Even successfully generating YAML output, Mistral and Gemma frequently introduce noisy information (88%-65% for Mistral with and without CoT, 98%-79% for Gemma) in the response (e.g., a key “Task" should have multiple values, Mistral generates multiple key-value pairs instead e.g., “Task_1:Training $\cdots$ Task_2: $\cdots$”), resulting in poor overall performance.
6 Mitigating Performance Format Bias: Actionable Recommendations
We propose methods as actionable recommendations to mitigate format bias in the performance of LLMs. Generally, three primary streams of techniques have been widely studied and applied to tackle LM biases: (1) Prompting Xu et al. (2024); Macedo et al. (2024); (2) Calibrating Roelofs et al. (2022); Li et al. (2024); and (3) Fine-tuning Schick et al. (2021); Ghaddar et al. (2021). While calibration techniques can only be used for white-box models, prompting and fine-tuning can be applied for both black-box (via API) and white-box ones. Therefore, we explore prompting and fine-tuning techniques to reduce format bias. We target mitigating the format bias of ChatGPT in Fig. 6, the strongest model that we benchmarked, on MMLU. We aim to reduce the wrapping bias (Section 5.2) due to resource limits, but our methods can be generalized to any model and format. We also verify (and confirm) our mitigation strategies on Gemma-2B-it, a medium-size model in Appx.-Tab. 18.
Figure 6: More demonstrations and repeating format instructions mitigate format bias. Finetuning mostly eliminates the format bias. The performance is reported using ChatGPT on MMLU (Appx.-Tab. 18 for num. results).
Demonstration(s) reduce(s) format bias.
As discussed in Section 5.1, LLMs show bias across formats possibly because of the token bias issue, causing LLMs to non-uniformly comprehend the format instructions. To address this, we examine whether demonstrations with formats can reduce such bias, as they are commonly utilized to enhance LLM’s comprehension of the task patterns Xie et al. (2022). Particularly, for each wrapping format in Section 5.2, we select $1$ and $5$ random samples from the auxiliary train data of MMLU and manually format the answers as demonstrations. The results are outlined in Fig. 6. Firstly, incorporating demonstrations typically enhances the FI scores of the model (from 72.99% to 79.19% and 85.31%) (i), with five demonstrations yielding the most. Secondly, we observe a notable decrease in the $BiasF_{o}$ score (ii) upon supplementing demonstrations. From (i), (ii) and Section 3.2, we conclude integrating demonstrations mitigates format bias.
Repeating format instructions reduces format bias.
We found that repeating instructions generally increases FI scores (i) across most formats except “Placeholder”, which can consequently lessen the mode’s token bias towards format instructions (Section 5.1). Using our two proposed criteria for effective format bias mitigation in Section 3.2, it is worth examining if this approach reduces $BiasF_{o}$, thereby being an effective mitigation. Our answer is yes. By repeating the wrapping instructions of ChatGPT thrice, we observed a decrease in the $BiasF_{o}$ (ii) score presented in Fig. 6. Combining (i) and (ii) suggests that this strategy is an effective mitigation. For "Placeholder," human investigation reveals that multiple placeholder instructions cause ChatGPT to be confused about where the placeholder is, making it frequently misunderstand and fail to follow this format instruction.
Fine-tuning with additional format data can eliminate format bias.
We hypothesize that completely solving the format token bias problem of LLMs necessitates finetuning them on format data so that they are familiar with tokens in format instructions evenly. We propose a simple data synthesis strategy for finetuning LLMs: we sample a small set of training data for all evaluated formats, with ratios inversely proportional to their systematic evaluation scores (Section 3.1). We chose $SysE$ scores over the $EstTrueE$ because they reflect the current model performance. Practically, based on ChatGPT’s zero-shot systematic performance on MMLU colored in blue in Appx.-Tab. 6, we approximate the formats’ performance ratios as “$1,1,\frac{1}{2},\frac{1}{2},\frac{1}{3},1,\frac{1}{3}$” from left-to-right, resulting in training data ratios of formats of “$1,1,2,2,3,1,3$”. We then preprocess the MMLU auxiliary training data according to these ratios, scaled by $500$ ($6500$ samples total), and train ChatGPT on this dataset. The finetuned results are plotted in Fig. 6. Firstly, after finetuning, the average FI score across all formats is nearly perfect at $99.58\%$ (ii). Secondly, the $BiasF_{o}$ score is significantly reduced from $235.33\%^{2}$ to $0.71\%^{2}$ (ii). These (i) and (ii) indicate finetuning largely eliminates format bias.
7 Conclusions
We introduce the pioneering systematic investigation of format bias in LLM performance, revealing significant biases across widely used formats for all models and benchmarks. Our method involves developing metrics to assess this bias and establishing criteria for effective mitigation. We then introduce prompting and fine-tuning techniques to alleviate format bias based on our evaluation findings. Our work aims to sharpen the focus of future LLM research toward fairer and more robust development.
Limitations
Our study has several limitations. Firstly, the metrics $EstTrue$ and $BiasF_{o}$ proposed in Section 3.1 and Section 3.2 are estimators, not exact measures. As discussed, determining $TrueE$ (Eq. 4) is infeasible, especially for large-scale experiments across various models and datasets. Achieving this would require extensive fine-tuning and comprehensive human evaluations, both prohibitively expensive and impractical in many scenarios. Our proposed metrics $EstTrue$ and $BiasF_{o}$ are handy for large-scale experiments with multiple models and datasets due to their fully automatic nature. We further propose Theorem 3.1 to validate the reliability of $TrueE$ statistically.
Secondly, our empirical evaluation of format bias is limited by computational and budget constraints to specific datasets, formats, and models. This restriction limits the generalizability of our findings and may obscure further insights that could be gained from expanding the experiments to include more formats, larger-scale datasets, and additional task categories.
Finally, while our study primarily attributes format bias to token bias in the training data of LLMs and proposes data-focused approach, it does not extensively explore other factors related to model architecture and training processes. This omission represents a significant area for future research, as more fundamental, architecture-level solutions could be crucial, for addressing format bias in LLMs. Our study underscores the importance of continued research dedicated to quantifying and mitigating format bias.
Ethical Considerations
Our work uncovers significant format bias in LLMs, raising concerns regarding fairness and potential discrimination in real-world applications.
Bias and fairness.
Format bias in LLMs can result in unfair treatment, especially in tasks where multiple possible formats can be used. Our research suggests ways to identify and mitigate format bias, aiming for fairer and more equitable LLM applications.
Societal impact.
Format bias in LLMs has the potential to disproportionately impact specific populations, as different demographics may have preferences for different communication formats. Further research is essential to fully understand its societal implications and ensure fairness across diverse demographics.
Acknowledgements
This research project is partially supported by the National Research Foundation Singapore under the AI Singapore Programme (AISG Award No: AISG2-TC-2023-010-SGIL), the Singapore Ministry of Education Academic Research Fund Tier 1 (Award No: T1 251RES2207), and the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-GC-2022-005). DXL and TS are supported by the A*STAR Computing and Information Science (ACIS) scholarship. We thank Goh Yisheng for his contribution in the initial stage of the project. We thank members of WING and Deep Learning Lab, NUS, as well as the anonymous reviewers for the constructive feedback.
Appendix A Format-Instruction Following Scorer
Algorithm 1 Format-Instruction Following Scorer
0:  Task $T$, language model $\mathcal{M}$, format constraints $C$, generated output $Y$.
0:  If $C$ includes wrapping characters, we denote as $\{W_{1},W_{2}\}$ and $is\_wrapping=True$.
0:  $output\_type$ is the data type required by $C$ when $T$ is not MCQ.
1:  if $is\_wrapping$ then
2:     return  False if (any of $\{W_{1},W_{2}\}\not\in Y$) or (number of $W_{1}\in Y$ + number of $W_{2}\in Y$neq 2).
3:     $ans$ = Extract string in between $\{W_{1},W_{2}\}$.
4:  else
5:     $ans=Y$
6:  end if
7:  if $T$ is MCQ then
8:     if MCQ output type is character identifier then
9:        return  True if $ans\in\{A,B,C,D\}$. False otherwise.
10:     else
11:        return  True if $ans\in$options’ values}. False otherwise.
12:     end if
13:  else
14:     return  True if we can parse $ans$ as an instance of the class $output\_type$. False otherwise.
15:  end if
Alg. 1 presents our heuristic algorithm for evaluating the format-instruction following capabilities of LLMs, which is used to compute $F_{C}$ in Eq. 1. The algorithm is divided into two three main parts:
  1. 1.
    Lines 1-6. These lines focus on examining the wrapping requirements by verifying the presence and correctness of the specified wrapping tokens.
  2. 2.
    Lines 7-12. These lines are dedicated to checking the formats of MCQ answers (Section 5.1).
  3. 3.
    Lines 13-15. These lines address the remaining formats, including list and mapping formats.
It is worth noting that Alg. 1 is highly adaptable; formats can be added or removed to tailor it for specific downstream applications.
Appendix B Theoretical Analysis: Reliability of $EstTrueE$
B.1 Proof of Theorem 3.1
Proof of Theorem 3.1.
We omit the case when $FI_{C}=0$ since in that case, we cannot estimate $TrueE$. By the definition in Theorem 3.1, we have $S_{C}$ generated answers that satisfy $C$. Let’s denote $k=S_{C}$ for simplicity. Let’s denote $k$ performance scores of answers satisfying $C$ as $x_{1},\cdots,x_{k}$, and $\bar{x}=\frac{\sum_{i=1}^{k}(x_{i})}{k}$ as the mean. Finally, $TrueE$ is the population mean of the performance scores, denoted as $\mu$.
Statement 1: $EstTrueE$ is consistent. From Eq. 5, by rewriting $EstTrueE$, we have $EstTrueE=\frac{1}{n}\cdot\sum_{i=1}^{k}(x_{i})\cdot\frac{n}{k}=\bar{x}$, which is an unbiased estimator of the average performance $TrueE$, i.e., $Bias(\bar{x})=0$ or $\lim_{k\to\infty}Bias(EstTrueE)=0$ (1). Now, let’s denote the variance of the performance scores as $\sigma^{2}$, then the variance of $EstTrueE$ is $Var(EstTrueE)=Var(\bar{x})=\frac{\sigma^{2}}{n}$ and $\lim_{k\to\infty}Var(EstTrueE)=0$ (2). From (1) and (2), by the Sufficient Condition for Consistency Amemiya (1985), we conclude that $EstTrueE$ is a consistent estimator.
Statement 2: $FI_{C}$ value. Let’s denote $s^{2}=\frac{1}{k-1}\sum_{i=1}^{k}(x_{i}-\bar{x})^{2}$ as the sample variance of the performance scores $x_{i}$s. It is well-known that $\frac{\sqrt{k}(\bar{x}-\mu)}{s}\sim t_{k-1}$. For estimating the population mean $\mu$ with finite population size $n$ and the type I error $\alpha$, we have the margin of error $\epsilon$:
$\epsilon\geq t_{\alpha/2,k-1}\cdot\sqrt{\frac{n-k}{n}\cdot\frac{s^{2}}{k}}$ (9)
where $\frac{n-k}{n}$ is the finite population correction factor. Eq. 9 is equivalent to:
$k\geq\frac{n-k}{n}\cdot\left(\frac{t_{\alpha/2,k-1}\cdot s}{\epsilon}\right)^{2}$ (10)
which yields
$k\geq\frac{1}{\frac{1}{n}+\left(\frac{\epsilon}{t_{\alpha/2,k-1}\cdot s}\right)^{2}}.$ (11)
then
$FI_{C}=\frac{k}{n}\geq\frac{1}{1+n\cdot\left(\frac{\epsilon}{t_{\alpha/2,k-1}\cdot s}\right)^{2}}.$ (12)
Statement 3: When $FI_{C}$ approaches $1$, $EstTrueE$ approaches $TrueE$. Since $EstTrueE$ by its definition in Eq. 5 is continuous with respect to $FI_{C}$ (Eq. 5), $S_{C}$ (Eq. 3) and $F_{C}$ (Eq. 3), therefore, we have the equality:
$\lim_{{FI_{C}}\to 100}{(EstTrueE)}=EstTrueE(FI_{C}=100)={TrueE}.$
B.2 Python Codes for Computing Reliability
1import numpy as np
2from scipy.stats import t
3import math
4
5def compute_sample_variance(data):
6 n = len(data)
7 mean = np.mean(data)
8 squared_deviations = [(x - mean) ** 2 for x in data]
9 sample_variance = sum(squared_deviations) / (n - 1)
10 return sample_variance
11
12def is_estimator_reliable(num_FI, list_eval_scores, num_samples=200):
13 ####### t-statistics #######
14 alpha = 0.05 # 5% significance level
15 df = num_FI - 1 # degrees of freedom
16 alpha_two_tailed = alpha / 2
17 t_statistic = t.ppf(1 - alpha_two_tailed, df)
18
19 #######Compute MOE_FI #######
20 epsilon = 0.05 # 5% margin of error
21 s = math.sqrt(compute_sample_variance(list_eval_scores))
22 return num_FI/num_samples > 1/(1 + num_samples * (epsilon/(t_statistic * s))**2)
Listing 1: Python codes for computing the reliability of $EstTrueE$ with margin of errors 5% performance with a significance level 5%.
Appendix C Detailed Discussions
We give the numerical results and discussions for all figures and points made in the main paper.
C.1 Multiple-choice Question (MCQ) Discussions
We evaluate Gemma, Mistral, and ChatGPT on the MMLU and BBH datasets using two prompting techniques, Zero-shot (ZS) and Zero-shot Chain-of-Thought (ZS-CoT) (Section 5.1). The prompts are specified in Section E.1. We report the $FI_{C},SysE,EstTrueE$ scores. The results are presented in Tab. 2. Additionally, Tab. 3, Tab. 4, and Tab. 5 are the distillation results of Tab. 2:
  1. 1.
    Tab. 3. For each model, we average its $EstTrueE$ performance overall benchmarks and prompting techniques. For each task, we average the $EstTrueE$ scores overall models and prompting techniques. The results of this table are plotted in Fig. 2 and discussed in Section 5.1.
  2. 2.
    Tab. 4. The purpose of this table is to compare the FI scores across formats. We average all the FI scores across models and tasks.
  3. 3.
    Tab. 5. The purpose of this table is to see whether CoT Wei et al. (2022) mitigates format bias. We average all the $EstTrueE$ scores over all models and benchmarks for each ZS and ZS-CoT prompting method.
MCQ type Char. Text.
MMLU
Gemma-7B-it (EstTrue-Acc) 0.53 / 27.25 8.10 / 18.63
Gemma-7B-it (Systematic-Acc) 0.12 / 10.32 0.17 / 4.86
Gemma-7B-it (FI) 22.47 / 37.87 2.10 / 26.09
Mistral-7B-it-v0.2 (EstTrue-Acc) 46.14 / 49.31 8.37 / 8.52
Mistral-7B-it-v0.2 (Systematic-Acc) 41.59 / 45.94 0.17 / 0.19
Mistral-7B-it-v0.2 (FI) 90.12 / 93.16 2.03 / 2.23
Llama-3.1-8B-it (EstTrue-Acc) 20.04 / 35.03 0.00 / 47.66
Llama-3.1-8B-it (Systematic-Acc) 1.79 / 27.16 0.00 / 29.05
Llama-3.1-8B-it (FI) 8.93 / 77.54 0.00 / 60.95
ChatGPT (EstTrue-Acc) 68.55 / 45.53 54.85 / 59.67
ChatGPT (Systematic-Acc) 66.20 / 42.22 12.71 / 26.31
ChatGPT (FI) 96.56 / 92.73 23.17 / 44.09
BBH
Gemma-7B-it (EstTrue-Acc) 42.11 / 23.05 0.00 / 15.11
Gemma-7B-it (Systematic-Acc) 0.40 / 13.00 0.00 / 6.80
Gemma-7B-it (FI) 0.95 / 56.40 0.00 / 45.00
Mistral-7B-it-v0.2 (EstTrue-Acc) 76.81 / 62.50 0.00 / 0.00
Mistral-7B-it-v0.2 (Systematic-Acc) 21.20 / 22.00 0.00 / 0.00
Mistral-7B-it-v0.2 (FI) 27.60 / 35.20 0.00 / 1.60
Llama-3.1-8B-it (EstTrue-Acc) 0.00 / 63.04 0.00 / 48.40
Llama-3.1-8B-it (Systematic-Acc) 0.00 / 34.80 0.00 / 36.40
Llama-3.1-8B-it (FI) 0.00 / 55.20 0.00 / 75.20
ChatGPT (EstTrue-Acc) 73.03 / 57.14 53.63 / 0.00
ChatGPT (Systematic-Acc) 26.00 / 16.0 53.20 / 0.00
ChatGPT (FI) 35.60 / 28.00 99.20 / 0.00
Table 2: MCQ results. Red results are unreliable results measured by Theorem 3.1 i.e., inequality Eq. 6 does not hold.
Char. Text. $BiasF_{o}$ (Var)
Models
Gemma 23.24 10.46 40.83
Mistral 58.69 4.22 741.74
Llama 5.01 48.53 473.56
ChatGPT 61.07 42.04 90.53
Tasks
MMLU 32.17 30.10 1.06
BBH 41.83 22.52 93.18
Table 3: Average. estimated true accuracy results of MCQ benchmarks across models and datasets (MMLU, BBH) with two MCQ formats (Character and Textual Value). This table is the distillation result of averaging all the estimated true performance scores across models (Models) and tasks (Tasks) in Tab. 3.
Format instruction-following bias.
The FI scores across formats are illustrated in Tab. 4. There is a notable difference between the scores for character-based and textual value-based formats. Among the models, ChatGPT follows the instructions best with FI score $52.42\%$. Below we present two examples of Gemma and Mistral failing to follow the format instructions:
Models Char. Text. Avg. $BiasF_{o}$ (Var)
Gemma 29.42 18.29 23.86 30.96
Mistral 61.52 1.46 31.49 901.80
Llama 2.23 67.22 34.73 1055.92
ChatGPT 63.22 41.61 52.41 116.74
Avg. 39.10 25.85 32.48 87.72
Table 4: MCQ format-instruction following scores across formats. This table is the distillation result of averaging all the FI scores across tasks and prompting techniques in Tab. 3.
Answer the following question by outputting the textual value of the choice without its character identifier. Wrap your final answer by <ANSWER> and </ANSWER>.
Question: Is the following sentence plausible? ‘‘John Carlson scored in the third period.’’
Answer choices: A. yes | B. no
Gemma:
<ANSWER>A</ANSWER>
Answer the following question by outputting the textual value of the choice without its character identifier. Wrap your final answer by <ANSWER> and </ANSWER>.
Question: Is the following sentence plausible? ‘‘Marcelo got on the end of a through ball.’’
Answer choices: A. yes | B. no
Mistral:
Based on the given sentence and the provided answer choices, the answer is: A. yes. The sentence describes Marcelo receiving a through ball and getting on the end of it, which is a common football (soccer) term for receiving a long, accurate pass and finishing a move with a shot or a goal.
ZS versus ZS-CoT.
We investigate the impact of Chain-of-Thought (CoT) Wei et al. (2022) on mitigating format bias. The average of estimated True Accuracy scores across models and tasks for the ZS prompting and ZS-CoT prompting are shown in Tab. 5. ZS prompting achieves a higher score in the character category (40.90%) compared to ZS-CoT prompting (33.10%). Similarly, for textual value format, ZS prompting scores higher than ZS-CoT prompting. However, the $BiasF_{o}$ is lower for the ZS-CoT model ($17.42\%^{2}$) compared to the ZS model ($42.42\%^{2}$), indicating that CoT slightly decreases the format bias.
Char. Text. $BiasF_{o}$
Zero-shot 40.90 27.88 42.42
Zero-shot Chain-of-Thought 33.10 24.75 17.42
Table 5: MCQ CoT versus non-CoT. This table is the distillation result of averaging all the Zero-shot and Zero-shot Chain-of-Thought scores across models and tasks in Tab. 3.
Reliability of the results.
From Tab. 2, we see that $21/32$ of the estimated $EstTrue$ results are reliable. The reliability of results in the MCQ output format varies across different models. Gemma-7B-it and Mistral-7B-it show significant unreliability in textual value format, evidenced by numerous red-marked scores due to models not following the format instructions to output correct formats. In contrast, ChatGPT’s results are significantly more reliable in the MMLU and BBH benchmarks ($7/8$), with only one unreliable result in the BBH textual format output.
C.2 Wrapping Discussions
We examine Gemma, Mistral, and ChatGPT on the MCQ datasets (MMLU,BBH) and generation datasets (GSM8K, HotpotQA, FairytaleQA) utilizing two prompting techniques, Zero-shot (ZS) and Zero-shot Chain-of-Thought (ZS-CoT) (Section 5.2). The prompts are also provided in Section E.2. We measure the $FI_{C},SysE,EstTrueE$. The results are shown in Tab. 6. Furthermore, Tab. 7, Tab. 8 and Tab. 9 are the distillation outcome of Tab. 6:
  1. 1.
    Tab. 7. For each model, we average its $EstTrueE$ performance overall benchmarks and prompting techniques. For each task, we average the $EstTrueE$ scores overall models and prompting techniques. This table is plotted in Fig. 3 and discussed in Section 5.2.
  2. 2.
    Tab. 8. The purpose of this table is to compare the FI scores across formats. We average all the FI scores across models and tasks.
  3. 3.
    Tab. 9. The purpose of this table is to see whether CoT Wei et al. (2022) mitigates format bias. We average all the $EstTrueE$ scores over all models and benchmarks for each ZS and ZS-CoT prompting method.
Wrapping type Special character Bolding Italicizing Brackets Parentheses Placeholder Quoting
MMLU
Gemma-7B-it (EstTrue-Acc) 35.59 / 20.28 41.28 / 44.27 49.85 / 74.18 36.36 / 32.95 36.68 / 20.12 46.45 / 25.77 60.41 / 74.06
Gemma-7B-it (Systematic-Acc) 27.82 / 20.28 21.66 / 17.73 26.64 / 27.89 28.55 / 27.28 10.53 / 12.96 29.80 / 21.96 2.64 / 2.37
Gemma-7B-it (FI) 78.16 / 100.00 52.47 / 39.60 53.44 / 37.60 78.52 / 82.80 28.71 / 64.40 64.15 / 85.20 4.37 / 3.20
Mistral-7B-it (EstTrue-Acc) 53.63 / 58.34 48.43 / 63.09 51.84 / 61.66 67.36 / 61.58 64.99 / 62.71 75.35 / 6.03 100.00 / 8.33
Mistral-7B-it (Systematic-Acc) 13.42 / 20.04 1.08 / 9.40 4.80 / 10.15 20.08 / 17.28 11.10 / 13.42 1.07 / 0.14 0.03 / 0.01
Mistral-7B-it (FI) 23.81 / 34.35 2.23 / 14.90 9.26 / 16.46 29.81 / 28.06 17.08 / 21.40 1.42 / 2.32 0.03 / 0.12
Llama-3.1-8B-it (EstTrue-F1) 52.06 / 55.95 25.08 / 54.42 48.54 / 89.11 92.36 / 19.63 39.28 / 32.00 65.18 / 68.12 34.67 / 42.83
Llama-3.1-8B-it (Systematic-F1) 38.05 / 37.92 12.93 / 28.82 20.48 / 32.57 11.61 / 8.02 31.80 / 30.39 56.53 / 27.91 26.89 / 16.91
Llama-3.1-8B-it (FI) 73.09 / 67.78 51.56 / 52.95 42.19 / 36.55 12.57 / 40.85 80.95 / 95.00 86.73 / 40.97 60.26 / 39.48
ChatGPT (EstTrue-Acc) 54.64/ 71.28 67.40 / 75.86 44.76 / 64.79 59.80 / 71.42 57.82 / 71.11 66.24 / 72.81 68.29 / 70.68
ChatGPT (Systematic-Acc) 48.54 / 63.64 66.59 / 48.59 38.24 / 36.77 31.65 / 60.86 28.54 / 60.57 63.88 / 50.09 26.72 / 30.26
ChatGPT (FI) 88.84 / 89.28 98.80 / 64.05 85.43 / 56.75 52.93 / 85.21 49.36 / 85.18 96.44 / 68.80 39.13 / 42.81
BBH
Gemma-7B-it (EstTrue-Acc) 25.00 / 16.00 49.09 / 38.38 52.94 / 24.47 63.04 / 47.34 36.73 / 26.09 7.07 / 3.76 60.00 / 20.00
Gemma-7B-it (Systematic-Acc) 24.00 / 16.00 21.60 / 15.20 10.80 / 9.20 23.20 / 19.60 14.40 / 16.80 5.20 / 3.20 2.40 / 0.40
Gemma-7B-it (FI) 96.00 / 100.00 44.00 / 39.60 20.40 / 37.60 36.80 / 41.40 39.20 / 64.40 73.60 / 85.20 4.00 / 2.00
Mistral-7B-it (EstTrue-Acc) 52.40 / 64.00 10.40 / 11.60 36.80 / 21.20 16.00 / 8.40 6.4 / 12.00 32.80 / 72.80 0.00 / 0.00
Mistral-7B-it (Systematic-Acc) 49.04 / 58.11 1.37 / 1.85 34.88 / 14.24 6.84 / 1.61 1.51 / 3.98 13.38 / 71.05 0.00 / 0.00
Mistral-7B-it (FI) 93.60 / 90.80 13.20 / 16.00 94.80 / 67.20 42.80 / 19.20 23.60 / 33.20 40.80 / 97.60 0.00 / 0.00
Llama-3.1-8B-it (EstTrue-F1) 57.00 / 51.41 14.03 / 45.14 9.90 / 34.74 42.86 / 27.97 50.40 / 29.32 57.14 / 66.32 49.64 / 21.00
Llama-3.1-8B-it (Systematic-F1) 48.80 / 36.40 7.20 / 26.00 4.00 / 13.20 12.00 / 8.00 50.40 / 29.20 52.80 / 25.20 27.20 / 18.40
Llama-3.1-8B-it (FI) 85.60 / 70.80 53.60 / 57.60 40.80 / 38.00 28.00 / 27.60 100.00 / 99.60 92.40 / 38.00 54.80 / 37.60
ChatGPT (EstTrue-Acc) 64.00 / 47.20 74.80 / 36.80 9.20 / 14.40 53.60 / 51.60 63.60 / 13.60 54.00 / 14.80 14.00 / 18.00
ChatGPT (Systematic-Acc) 64.00 / 16.80 74.80 / 30.62 9.20 / 10.02 51.67 / 38.60 57.24 / 3.75 54.00 / 14.80 3.19 / 0.58
ChatGPT (FI) 100.00 / 35.60 100.00 / 83.20 100.00 / 69.60 96.40 / 74.80 90.00 / 27.60 100.00 / 100.00 22.80 / 3.20
GSM8K
Gemma-7B-it (EstTrue-$F1$) 3.65 / 5.00 0.99 / 3.13 5.20 / 1.46 7.45 / 0.42 0.00 / 0.00 9.13 / 9.92 0.0 / 0.0
Gemma-7B-it (Systematic-$F1$) 2.54 / 2.45 0.50 / 2.00 4.26 / 1.19 3.50 / 0.17 0.00 / 0.00 4.52 / 4.71 0.0 / 0.0
Gemma-7B-it (FI) 69.50 / 49.00 50.50 / 64.00 82.00 / 81.50 47.00 / 40.05 2.50 / 0.50 49.50 / 47.50 0.0 / 0.0
Mistral-7B-it (EstTrue-F1) 4.03 / 25.74 9.03 / 31.61 2.87 / 30.76 2.57 / 46.98 1.29 / 39.44 3.28 / 39.37 0.00 / 73.52
Mistral-7B-it (Systematic-F1) 3.43 / 23.43 1.40 / 4.11 1.42 / 20.76 1.67 / 38.76 0.60 / 24.26 3.28 / 38.78 0.00 / 6.25
Mistral-7B-it (FI) 85.00 / 91.00 15.50 / 13.00 49.50 / 67.50 65.00 / 82.50 46.50 / 61.50 100.00 / 98.50 5.00 / 8.50
Llama-3.1-8B-it (EstTrue-F1) 48.22 / 50.76 12.90 / 26.24 8.13 / 13.17 7.49 / 49.59 6.67 / 70.50 7.50 / 58.29 0.00 / 0.00
Llama-3.1-8B-it (Systematic-F1) 47.50 / 50.00 10.00 / 18.50 5.00 / 11.00 7.00 / 30.50 5.50 / 49.00 7.50 / 58.00 0.00 / 0.00
Llama-3.1-8B-it (FI) 98.50 / 98.50 77.50 / 70.50 61.50 / 83.50 93.50 / 61.50 82.50 / 69.50 100.00 / 99.50 0.00 / 0.00
ChatGPT (EstTrue-F1) 19.54 / 43.98 22.95 / 24.36 21.22 / 30.57 21.27 / 69.00 22.02 / 63.83 23.03 / 60.25 16.43 / 24.01
ChatGPT (Systematic-F1) 19.44 / 43.98 22.84 / 23.39 21.12 / 24.15 20.74 / 67.62 21.25 / 62.24 23.03 / 59.05 9.78 / 14.65
ChatGPT (FI) 99.50 / 100.00 99.50 / 96.00 99.50 / 79.00 97.50 / 98.50 96.50 / 97.50 100.00 / 98.00 59.50 / 61.00
HotpotQA
Gemma-7B-it (EstTrue-$F1$) 14.12 / 9.88 21.43 / 32.11 19.83 / 27.06 23.63 / 30.44 0.00 / 0.00 43.70 / 53.62 2.33 / 6.60
Gemma-7B-it (Systematic-$F1$) 4.59 / 5.53 9.00 / 12.20 7.93 / 8.93 3.90 / 14.00 0.00 / 0.00 5.90 / 9.92 0.03 / 0.03
Gemma-7B-it (FI) 32.50 / 56.00 42.00 / 38.00 40.00 / 33.00 16.50 / 46.00 3.50 / 2.50 13.50 / 18.50 1.50 / 0.50
Mistral-7B-it (EstTrue-F1) 12.86 / 11.43 25.84 / 29.21 20.93 / 14.56 16.93 / 13.20 15.39 / 13.21 20.41 / 21.58 0.00 / 25.00
Mistral-7B-it (Systematic-F1) 7.27 / 3.83 8.27 / 3.36 6.91 / 4.95 16.51 / 10.76 14.55 / 10.24 19.70 / 14.75 0.00 / 0.05
Mistral-7B-it (FI) 56.50 / 33.50 32.00 / 11.50 33.00 / 34.00 97.50 / 81.50 94.50 / 77.50 96.50 / 91.50 0.00 / 0.20
Llama-3.1-8B-it (EstTrue-F1) 20.63 / 21.65 21.93 / 14.11 5.34 / 5.92 19.17 / 18.08 19.79 / 17.59 20.00 / 21.39 0.00 / 0.00
Llama-3.1-8B-it (Systematic-F1) 13.00 / 10.50 21.50 / 11.50 5.00 / 4.00 18.50 / 8.50 19.50 / 9.50 20.00 / 20.00 0.00 / 0.00
Llama-3.1-8B-it (FI) 63.00 / 48.50 98.00 / 81.50 93.50 / 67.50 96.50 / 47.00 98.50 / 54.00 100.00 / 93.50 0.00 / 0.00
ChatGPT (EstTrue-F1) 29.86 / 27.52 41.00 / 33.14 35.39 / 28.96 23.94 / 35.48 29.30 / 34.83 38.72 / 28.69 41.52 / 16.97
ChatGPT (Systematic-F1) 25.24 / 27.11 40.59 / 30.82 33.45 / 26.64 17.00 / 33.36 23.46 / 33.44 38.72 / 27.69 11.73 / 7.13
ChatGPT (FI) 84.50 / 98.50 99.00 / 93.00 94.50 / 92.00 71.50 / 94.00 80.05 / 96.00 100.00 / 96.50 28.50 / 42.00
FairytaleQA
Gemma-7B-it (EstTrue-$F1$) 17.42 / 29.72 8.91 / 0.97 8.12 / 14.50 22.13 / 18.62 0.00 / 0.00 20.64 / 22.05 0.00 / 0.00
Gemma-7B-it (Systematic-$F1$) 6.62 / 11.74 4.68 / 0.64 4.75 / 9.79 1.77 / 1.21 0.00 / 0.00 2.58 / 4.08 0.00 / 0.00
Gemma-7B-it (FI) 38.00 / 39.50 52.50 / 66.00 58.50 / 67.50 8.00 / 6.50 0.00 / 0.00 12.50 / 18.50 0.00 / 0.00
Mistral-7B-it (EstTrue-F1) 27.19 / 22.20 23.78 / 50.00 47.36 / 29.49 32.42 / 25.90 30.33 / 22.46 36.07 / 31.77 19.50 / 20.00
Mistral-7B-it (Systematic-F1) 22.16 / 18.54 3.21 / 0.50 18.47 / 15.19 32.42 / 25.00 29.73 / 21.00 35.89 / 31.62 0.39 / 1.30
Mistral-7B-it (FI) 81.50 / 83.50 13.50 / 1.00 39.00 / 51.50 100.00 / 96.50 98.00 / 93.50 99.50 / 99.50 2.00 / 6.50
Llama-3.1-8B-it (EstTrue-F1) 49.62 / 36.78 48.35 / 39.96 45.42 / 35.75 48.85 / 40.76 48.63 / 42.48 52.55 / 40.39 0.00 / 0.00
Llama-3.1-8B-it (Systematic-F1) 48.13 / 36.41 48.35 / 36.76 45.42 / 29.85 48.85 / 20.79 48.39 / 26.55 52.29 / 39.18 0.00 / 0.00
Llama-3.1-8B-it (FI) 97.00 / 99.00 100.00 / 92.00 100.00 / 83.50 100.00 / 51.00 99.50 / 62.50 99.50 / 97.00 0.00 / 0.00
ChatGPT (EstTrue-F1) 41.93 / 31.95 46.08 / 32.84 48.11 / 33.46 41.53 / 38.25 38.25 / 34.82 46.83 / 32.85 45.78 / 27.75
ChatGPT (Systematic-F1) 38.58 / 31.47 46.08 / 31.86 48.11 / 31.96 41.33 / 38.06 45.91 / 34.30 46.83 / 32.85 27.24 / 14.71
ChatGPT (FI) 92.00 / 98.50 100.00 / 97.00 100.00 / 95.50 99.50 / 99.50 99.50 / 98.50 100.00 / 100.00 59.50 / 53.00
Table 6: Wrapping results. Red results are unreliable results measured by Theorem 3.1 i.e., inequality Eq. 6 does not hold.
Special Character Bolding Italicizing Brackets Parentheses Placeholder Quoting $BiasF_{o}$ (Var)
Models
Gemma 17.69 22.23 28.24 23.35 9.86 25.88 22.25 31.31
Mistral 32.87 29.33 31.05 28.46 26.02 33.23 24.64 9.20
LLama 44.41 30.16 28.98 36.78 35.67 45.69 18.85 74.86
ChatGPT 40.15 44.91 32.65 45.24 42.72 43.82 32.57 26.03
Average 33.78 31.65 30.23 33.46 28.57 37.15 24.58 14.14
Tasks
MMLU 50.56 52.48 60.59 55.18 48.09 53.24 58.65 16.58
BBH 47.13 34.96 24.68 38.98 29.77 38.59 26.32 54.58
GSM8K 25.12 16.40 14.17 25.55 25.47 26.35 14.25 28.51
HotpotQA 18.50 27.35 19.75 22.59 16.27 29.41 11.39 33.66
FairytaleQA 32.10 31.36 32.78 30.79 28.11 35.40 14.13 42.16
Average 34.68 32.51 30.39 34.62 29.54 36.60 24.95 13.39
Table 7: Avg. estimated true accuracy results of benchmarks across models and datasets with seven Wrapping formats (Special Character, Bolding, Italicizing, Brackets, Parentheses, Placeholder, Quoting). This table is the distillation result of averaging all the estimated true performance scores across models and benchmarks in Tab. 6.
Format instruction-following bias.
The FI scores over formats are provided in Tab. 8. Overall, LLMs exhibit significant format-following bias across formats with a variance of FI scores of $345.85\%^{2}$. Among the models, ChatGPT follows the instructions best with an average FI Score of $85.01\%$. The “Special Character” wrapping format has the highest FI score of $75.05\%$. Following it is the “Placeholder” wrapping format also shows a high FI score of $72.47\%$, suggesting it is another effective format for ensuring instruction adherence. In contrast, the “Quoting” wrapping format has the lowest FI score of $18.55\%$. This significant drop compared to other formats suggests that quoting is the least effective method for wrapping instructions, possibly causing confusion or misinterpretation by the models. Below we present two examples of Gemma and Mistral failing to follow the format instructions:
Model Special Character Bolding Italicizing Brackets Parentheses Placeholder Quoting Avg. $BiasF_{o}$ (Var)
Gemma 62.27 46.07 49.51 37.06 15.33 45.58 1.36 36.74 384.31
Mistral 63.00 15.36 40.94 67.53 54.72 64.48 4.04 44.30 553.55
LLama 80.18 73.52 64.70 55.85 84.21 84.76 23.01 66.60 413.00
ChatGPT 94.77 93.49 88.35 88.93 88.69 95.06 45.79 85.01 263.71
Avg. 75.05 57.11 59.60 60.87 60.74 72.47 18.55 57.76 345.85
Table 8: Avg. Following Instruction Score over all the wrapping formats. This table is the distillation result of averaging all the FI scores across models and benchmarks in Tab. 6.
Answer the following question without any explanation. Wrap your final answer using triple quotation marks.
Question (HotpotQA): ‘‘What was the MGM Grand Garden Arena in which Britney Spears recorded fourth video album originally known as?’’
Context:...
Gemma: MGM Grand Garden Special Events
Answer the following question without any explanation. Wrap your final answer using triple quotation marks.
Question (BBH): Is the following sentence plausible? ‘‘David Silva took a throw-in.’’
Mistral: A. ‘‘yes’’
ZS versus ZS-CoT.
The average of estimated True Accuracy scores across models and tasks for the ZS and ZS-CoT prompting are shown in Tab. 9. For the majority of the wrapping methods (“Special Character”, “Bolding”, “Italicizing”, and “Brackets”), the ZS-CoT model generally shows higher or comparable performance to the ZS model. The “Italicizing” shows a significant improvement when using ZS-CoT, with a jump from 28.28% (ZS) to 32.19% (ZS-CoT). However, the $BiasF_{o}$ metric shows a considerable difference between $11.39\%^{2}$ (ZS) and $16.22\%^{2}$ (ZS-CoT). This depicts that while applying CoT may improve accuracy in certain methods, it does not generally reduce format bias.
Special Character Bolding Italicizing Brackets Parentheses Placeholder Quoting $BiasF_{o}$
Zero-shot 34.31 30.66 28.28 33.82 28.77 35.93 26.09 11.39
Zero-shot Chain-of-Thought 33.25 32.65 32.19 33.09 28.36 38.37 24.46 16.22
Table 9: Avg.Estimated Accuracy of non CoT versus CoT for wrapping methods. This table is the distillation result of averaging all the Zero-shot and Zero-shot Chain-of-Thought scores across models and tasks in Tab. 6.
Reliability of the results.
Overall, $80\%$ of the $EstTrue$ results are reliable. Gemma-7B-it shows mixed reliability, with some red-marked scores indicating unreliable results, particularly in the “Quoting” format. This is because Gemma failed to follow the quoting instruction to quote the final answer. Mistral-7B-it exhibits similar variability, with some unreliable scores in “Quoting” and “Placeholder” formats. ChatGPT generally demonstrates mostly reliable results, with only 1 quoting result unreliable.
C.3 List Discussions
Listing type Python Bullet Spe. Char. Newline
SciDocsRR
Gemma-7B-it (EstTrue-$mAP$) 0.0 / 61.65 0.0 / 73.0 0.0 / 60.00 0.0 / 60.15
Gemma-7B-it (Systematic-$mAP$) 0.0 / 15.72 0.0 / 1.46 0.0 / 0.90 0.0 / 28.27
Gemma-7B-it (FI) 0.0 / 25.50 0.0 / 2.00 0.0 / 1.50 0.0 / 47.00
Mistral (EstTrue-$mAP$) 50.21 / 52.61 0.00 / 0.00 0.00 / 0.00 78.08 / 58.36
Mistral (Systematic-$mAP$) 37.41 / 9.47 0.00 / 0.00 0.00 / 0.00 18.35 / 27.14
Mistral (FI) 74.50 / 18.00 0.00 / 0.00 0.00 / 0.00 23.50 / 46.50
Llama-3.1-8B-it (EstTrue-$F1$) 32.10 / 0.00 38.80 / 0.00 0.00 / 0.00 35.78 / 0.00
Llama-3.1-8B-it (Systematic-$F1$) 6.26 / 0.00 3.88 / 0.00 0.00 / 0.00 35.78 / 0.00
Llama-3.1-8B-it (FI) 19.50 / 0.00 10.00 / 0.00 0.00 / 0.00 100.00 / 0.00
ChatGPT (EstTrue-$mAP$) 35.29 / 50.17 49.94 / 59.64 55.69 / 57.78 38.54 / 57.56
ChatGPT (Systematic-$mAP$) 33.17 / 28.60 49.19 / 25.05 55.69 / 37.85 35.46 / 35.41
ChatGPT (FI) 94.00 / 57.00 98.50 / 42.00 100.00 / 65.50 92.00 / 61.50
SemEval2017
Gemma-7B-it (EstTrue-$F1$) 4.00 / 8.86 7.10 / 7.20 4.80 / 13.50 7.21 / 3.25
Gemma-7B-it (Systematic-$F1$) 0.04 / 1.64 1.80 / 2.10 4.80 / 13.50 7.21 / 1.51
Gemma-7B-it (FI) 1.00 / 18.50 25.50 / 29.15 100.00 / 100.00 100.00 / 46.50
Mistral (EstTrue-$F1$) 34.82 / 30.24 23.2 / 0.00 0.00 / 13.57 12.17 / 20.84
Mistral (Systematic-$F1$) 33.95 / 24.19 23.20 / 0.00 0.00 / 10.72 12.17 / 20.84
Mistral (FI) 97.50 / 80.00 100.00 / 100.00 0.00 / 79.00 100.00 / 100.00
Llama-3.1-8B-it (EstTrue-$F1$) 34.82 / 0.00 0.22 / 0.00 0.00 / 0.00 0.00 / 10.17
Llama-3.1-8B-it (Systematic-$F1$) 33.95 / 0.00 0.22 / 0.00 0.00 / 0.00 0.00 / 10.17
Llama-3.1-8B-it (FI) 97.50 / 0.00 100.00 / 100.00 0.00 / 75.50 100.00 / 100.00
ChatGPT (EstTrue-$F1$) 42.25 / 15.33 8.87 / 16.46 32.19 / 16.33 37.16 / 22.87
ChatGPT (Systematic-$F1$) 39.51 / 6.04 8.87 / 16.13 31.07 / 15.51 37.16 / 22.75
ChatGPT (FI) 93.50 / 39.39 100.00 / 97.97 96.50 / 94.94 100.00 / 99.49
Table 10: List results. Red results are unreliable results measured by Theorem 3.1 i.e., inequality Eq. 6 does not hold.
We assess Gemma, Mistral, and ChatGPT with two prompting techniques, Zero-shot (ZS) and Zero-shot Chain-of-Thought (ZS-CoT) (Section 5.3) on two benchmarks SciDocsRR and SemEval2017. Our prompts are provided in E.3. We utilize $FI_{C},SysE,TrueE$ as our evaluation metrics. The results are illustrated in Tab. 10. In addition, Tab. 11, Tab. 12 and Tab. 13 are the distillation results of Tab. 10:
  1. 1.
    Tab. 11. For each model, we average its $EstTrueE$ performance overall benchmarks and prompting techniques. For each task, we average the $EstTrueE$ scores overall models and prompting techniques. This table is drawn in Fig. 4 and its discussions are conducted in Section 5.3.
  2. 2.
    Tab. 12. The purpose of this table is to compare the FI scores across formats. We average all the FI scores across models and tasks.
  3. 3.
    Tab. 13. The purpose of this table is to see whether CoT Wei et al. (2022) mitigates format bias. We average all the $EstTrueE$ scores over all models and benchmarks for each ZS and ZS-CoT prompting method.
Python Bullet Special Character Newline $BiasF_{o}$ (Var)
Models
Gemma 17.12 18.25 15.12 16.21 1.32
Mistral 41.98 5.80 3.39 42.37 353.80
LLama 16.73 9.76 0.00 11.49 36.64
ChatGPT 35.76 33.73 40.50 39.03 7.08
Average 27.90 16.88 14.75 27.27 46.97
Tasks
SemEval2017 20.54 6.09 7.82 13.49 31.86
SciDocsRR 35.25 27.67 21.68 41.06 54.12
Average 27.90 16.88 14.75 27.27 46.97
Table 11: Avg estimated true F1 (SemEval2017) and MAP (SciDocsRR) scores over benchmarks (SemEval2017, SciDocsRR) and models (Gemma-7B-it, Mistral, ChatGPT) across 4 widely used list formats (Python, Bullet, Special Character, Newline)
Format instruction-following bias.
From Tab. 12, the “Newline list” format has the highest FI score of $68.29\%$, while the FI scores of other formats are around $40-50\%$. Among the models, ChatGPT follows the instructions best with FI score $83.27\%$. Specifically, ChatGPT excels at following instructions for all formats. In comparison, Mistral performs poorly with “Special character list”, while Gemma struggles the most with “Python” and “Bullet point list”. Explaining these phenomena, we observe that Gemma often misinterprets the task of converting textual samples into lists of numbers as a coding task, resulting in answers that do not match the desired format or content. We take an example in dataset SciDocsRR Section D.1 with detailed prompt as in Section E.3
Given a query, and a list of documents.
Query: Declarative Terrain Modeling for Military Training Games
Documents: [...] You are required to transform the list of documents into a binary list of 1 or 0 where 1 indicates the document brings useful information to the topic, and 0 indicates the document does not bring useful information to the topic. Generate your binary list as a Python list Gemma: You should not use any packages such as sklearn, gensim, and NLTK. You are required to write a short explanation for your binary list.
Given a query, and a list of documents.
Query: Declarative Terrain Modeling for Military Training Games
Documents: [...] You are required to transform the list of documents into a binary list of 1 or 0 where 1 indicates the document brings useful information to the topic, and 0 indicates the document does not bring useful information to the topic. Generate your binary list using bullet points Gemma: Your binary list must be in the following format: [1, 0, 1, 0, 0, 0, 1, 1, 0, 1, 0, 1, 1, 1, 1, 0, 1, 1, 1, 0].
Model Python Bullet Special Character Newline Avg. $BiasF_{o}$ (Var)
Gemma 7.51 13.41 34.55 42.40 24.46 277.69
Mistral 67.50 50.00 19.75 67.50 51.19 507.31
ChatGPT 70.97 84.61 89.24 88.25 83.27 71.13
LLama 29.25 52.50 18.88 75 43.91 470.51
Avg. 43.81 50.13 40.60 68.29 50.71 153.04
Table 12: Avg Following Instruction scores over benchmarks (SemEval2017, SciDocsRR) and models (Gemma-7B-it, Mistral, ChatGPT) across 4 widely used list formats (Python, Bullet, Special Character, Newline). This table is the distillation result of averaging all the FI scores across models and benchmarks in Tab. 10.
ZS versus ZS-CoT.
The results, detailed in Tab. 13 indicate that prompting with ZS-CoT substantially enhances model performance across various formats. Moreover, ZS-CoT effectively reduces format bias, as evidenced by the $BiasF_{o}$ metric decreasing from about $54\%^{2}$ to $22\%^{2}$. From this, we conclude that CoT reduces format bias.
Python Bullet Special Character Newline $BiasF_{o}$ (Var)
Zero-shot 29.19 15.13 11.05 25.45 54.48
Zero-shot Chain-of-Thought 26.61 18.64 18.46 29.09 22.40
Table 13: Avg estimated true F1 (SemEval2017) and MAP (SciDocsRR) scores of non-CoT versus CoT for list formats. This table is the distillation result of averaging all the scores across models and benchmarks in Tab. 10.
Reliability of the results.
From Tab. 10, $67\%$ of the $EstTrue$ results are reliable. However, some scores of Gemma-7B-it and LLama on these benchmarks are red-marked, indicating unreliable results of this model. In contrast, the ChatGPT’s results are perfectly reliable.
C.4 Mapping Discussions
Mapping type JSON YAML
SciREX Easy
Gemma-7B-it (EstTrue-F1) 14.60 / 20.84 18.20 / 0.82
Gemma-7B-it (Systematic) 3.54 / 3.79 3.03 / 0.10
Gemma-7B-it (FI) 24.24 / 18.18 16.64 / 12.12
Mistral-7B-it (EstTrue-F1) 28.83 / 32.82 0.00 / 0.00
Mistral-7B-it (Systematic) 11.36 / 32.33 0.00 / 0.00
Mistral-7B-it (FI) 39.39 / 98.48 0.00 / 3.03
Llama-3.1-8B-it (EstTrue-F1) 25.85 / 22.86 19.07 / 28.57
Llama-3.1-8B-it (Systematic) 2.35 / 14.55 6.36 / 15.15
Llama-3.1-8B-it (FI) 9.09 / 63.63 33.34 / 53.03
ChatGPT (EstTrue-F1) 35.99 / 22.40 23.63 / 26.60
ChatGPT (Systematic) 32.72 / 19.69 22.92 / 20.15
ChatGPT (FI) 90.90 / 87.87 96.96 / 75.75
SciREX Medium
Gemma-7B-it (EstTrue-F1) 18.17 / 5.27 0.00 / 1.87
Gemma-7B-it (Systematic) 3.03 / 0.88 0.00 / 0.17
Gemma-7B-it (FI) 16.67 / 16.67 18.18 / 9.09
Mistral-7B-it (EstTrue-F1) 26.48 / 23.81 18.97 / 20.83
Mistral-7B-it (Systematic) 21.27 / 23.81 1.15 / 0.25
Mistral-7B-it (FI) 80.30 / 100.00 6.06 / 1.20
Llama-3.1-8B-it (EstTrue-F1) 40.80 / 35.01 31.64 / 27.10
Llama-3.1-8B-it (Systematic) 9.89 / 28.12 25.41 / 21.76
Llama-3.1-8B-it (FI) 24.24 / 80.30 80.30 / 80.30
ChatGPT (EstTrue-F1) 29.07 / 27.29 36.55 / 22.70
ChatGPT (Systematic) 28.19 / 26.47 21.60 / 22.70
ChatGPT (FI) 96.96 / 96.96 59.09 / 100.00
SciREX Hard
Gemma-7B-it (EstTrue-F1) 34.40 / 29.18 1.65 / 0.87
Gemma-7B-it (Systematic) 4.17 / 10.61 0.25 / 0.04
Gemma-7B-it (FI) 12.12 / 36.36 15.15 / 4.55
Mistral-7B-it (EstTrue-F1) 22.44 / 30.34 12.54 / 15.95
Mistral-7B-it (Systematic) 20.40 / 26.66 1.71 / 1.58
Mistral-7B-it (FI) 90.90 / 87.87 13.63 / 9.90
Llama-3.1-8B-it (EstTrue-F1) 39.33 / 39.66 27.39 / 26.05
Llama-3.1-8B-it (Systematic) 5.36 / 33.05 25.32 / 24.08
Llama-3.1-8B-it (FI) 13.63 / 92.42 83.33 / 92.42
ChatGPT (EstTrue-F1) 20.25 / 22.57 11.76 / 12.07
ChatGPT (Systematic) 19.64 / 22.23 11.59 / 10.43
ChatGPT (FI) 96.96 / 98.48 98.48 / 86.36
Table 14: Mapping results. Red results are unreliable results measured by Theorem 3.1 i.e., inequality Eq. 6 does not hold.
JSON YAML Average $BiasF_{o}$ (Var)
Models
Gemma 20.42 3.91 12.17 68.14
Mistral 27.46 11.39 19.43 64.56
Llama 33.92 26.64 30.28 13.25
ChatGPT 26.27 22.22 24.25 4.10
Tasks
Easy 25.53 14.61 20.07 29.79
Medium 25.74 19.96 22.85 8.36
Hard 29.77 13.54 21.66 65.87
Table 15: Avg estimated true F1 scores over benchmarks (SciREX Easy, SciREX Medium and SciREX Hard) and models (Gemma-7B-it, Mistral, ChatGPT) across 2 widely used mapping formats (JSON and YAML). This table is the distillation result of averaging all the estimated true performance scores across models and benchmarks in Tab. 14.
We select Gemma, Mistral, and ChatGPT for our evaluation, using two prompting techniques: Zero-shot (ZS) and Zero-shot Chain-of-Thought (ZS-CoT) (Section 5.4). These models are tested on the SciREX dataset across three difficulty levels: Easy, Medium, and Hard. Detailed prompt specifications are provided in Section E.4. We calculate $FI_{C},SysE,EstTrueE$ with the results presented in Tab. 14. Furthermore, Tab. 15, Tab. 16 and Tab. 17 are distilled from Tab. 14:
  1. 1.
    Tab. 15. For each model, we average its $EstTrueE$ performance overall benchmarks and prompting techniques. For each task, we average the $EstTrueE$ scores overall models and prompting techniques. This table is numerical results of Fig. 5 and discussed in Section 5.4.
  2. 2.
    Tab. 16. The purpose of this table is to compare the FI scores across formats. We average all the FI scores across models and tasks.
  3. 3.
    Tab. 17. The purpose of this table is to see whether CoT Wei et al. (2022) mitigates format bias. We average all the $EstTrueE$ scores over all models and benchmarks for each ZS and ZS-CoT prompting method.
Format instruction-following bias.
From this Tab. 16, noticeably, there is a significant difference in the FI scores of two mapping formats (“JSON” and “YAML”). The average FI scores of the “JSON” mapping format is $60.98\%$, which is much higher than “YAML” mapping format ($43.96\%$). Overall, among models, ChatGPT yields the best result of $90.40\%$. Mistral fails to follow “YAML” format instruction, e.g. for the question in dataset SciREX Easy Section D.1 with detailed promt as in Section E.4
Extract the entities reflecting the tasks in the following document:
Document: Published as a conference paper at ICLR 2017 QUERY - REDUCTION NETWORKS FOR QUESTION ANSWERING.
Your output must be in YAML format: ‘‘‘‘‘‘ Task: [...] ’’’’’’. Wrap your final answer by <ANSWER> and </ANSWER>.
Mistral: ‘‘‘yaml
- TASK_1: Prepare conference paper
- TASK_2: Submit conference paper to ICLR 2017
- TASK_3: Research and develop Query-Reduction Networks for Question Answering
<ANSWER>
‘‘‘yaml
- TASK_1: Prepare conference paper
- TASK_2: Submit conference paper to ICLR 2017
- TASK_3: Research and develop Query-Reduction Networks for Question Answering’’’
</ANSWER>
Model JSON YAML Avg. $BiasF_{o}$
Gemma 20.71 12.62 16.66 16.34
Mistral 82.82 5.13 43.98 1509.00
Llama 45.70 71.97 58.84 172.46
ChatGPT 94.69 86.11 90.40 18.41
Avg. 60.98 43.96 52.47 72.45
Table 16: Avg FI scores over benchmarks and models across 2 widely used mapping formats (JSON and YAML). This table is the distillation result of averaging all the FI scores across models and benchmarks in Tab. 14.
ZS versus ZS-CoT.
JSON YAML $BiasF_{o}$
Zero-shot 28.02 16.79 31.55
Zero-shot Chain-of-Thought 26.01 15.29 28.73
Table 17: Avg ZS and ZS-CoT scores over benchmarks and models across 2 widely used mapping formats (JSON and YAML). This table is the distillation results across models and benchmarks in Tab. 14.
From Tab. 17, it is evident that the performance of ZS prompting surpasses that of ZS-CoT for both formats. Upon comparing the $BiasF_{o}$ across prompting techniques, we conclude that CoT Wei et al. (2022) does not mitigate format bias.
Reliability of the results.
From Tab. 14, $77\%$ of the $EstTrue$ results are reliable. The reliability of the results in the mapping output format shows variability across different models and formats. Noticeably, “YAML” mapping format results are less reliable than “JSON” ones. On the other hand, ChatGPT illustrates its high reliability in all mapping formats while Mistral-7B-it and Gemma-7B-it are opposite, and all the results in the “YAML” mapping format of these models are unreliable.
C.5 Mitigating Format Bias Results
Index Wrapping type Special character Bolding Italicizing Brackets Parentheses Placeholder Quoting Avg. $BiasF_{o}$ (Var)
No demo (Zero-shot)
1 ChatGPT (EstTrue-Acc) 54.63 67.39 44.76 59.79 57.82 66.23 68.28 235.33
2 ChatGPT (Systematic) 48.54 66.59 38.24 31.65 28.54 63.88 26.72 532.75
3 ChatGPT (FI) 88.84 98.80 85.43 52.93 49.36 96.44 39.13 72.99 61.12
Repeat format prompt thrice
4 ChatGPT (EstTrue-Acc) 60.09 67.88 55.65 61.99 63.71 30.31 68.28 146.79
5 ChatGPT (Systematic) 56.65 66.98 49.93 35.74 51.63 2.85 33.13 377.66
6 ChatGPT (FI) 94.26 98.67 89.71 57.65 81.03 9.40 48.52 68.46 884.34
1 demo
7 ChatGPT (EstTrue-Acc) 55.12 65.08 47.18 52.23 56.13 65.92 63.60 172.69
8 ChatGPT (Systematic) 50.54 64.49 43.98 40.02 31.02 62.19 28.10 397.62
9 ChatGPT (FI) 91.68 99.09 93.22 76.61 55.26 94.34 44.18 79.20 43.75
5 demos
10 ChatGPT (EstTrue-Acc) 51.77 58.30 45.21 46.79 52.52 62.84 55.24 111.78
11 ChatGPT (Systematic) 51.18 56.66 40.69 41.36 39.78 60.88 27.72 259.37
12 ChatGPT (FI) 98.85 97.19 90.01 88.39 75.74 96.88 50.18 85.32 32.93
Finetuned
13 ChatGPT (EstTrue-Acc) 74.02 74.73 71.53 73.88 74.09 74.27 74.19 0.71
14 ChatGPT (Systematic) 73.99 74.11 71.52 73.66 73.47 74.15 73.70 0.11
15 ChatGPT (FI) 99.96 99.17 99.98 99.69 99.16 99.83 99.33 99.59 0.93
Index Wrapping type Special character Bolding Italicizing Brackets Parentheses Placeholder Quoting Avg. $BiasF_{o}$ (Var)
No demo (Zero-shot)
16 Gemma-2B (EstTrue-Acc) 30.64 6.88 1.90 0.0 0.0 8.41 0.0 104.81
17 Gemma-2B (Systematic) 23.26 4.32 0.78 0.0 0.0 8.32 0.0 63.00
18 Gemma-2B (FI) 76.24 62.83 41.06 0.0 0.0 98.91 0.0 39.86 1443.70
Repeat format prompt quintice
19 Gemma-2B (EstTrue-Acc) 58.40 0.61 32.26 0.00 0.00 55.52 0.00 636.50
20 Gemma-2B (Systematic) 56.00 0.40 14.00 0.00 0.00 54.00 0.00 576.77
21 Gemma-2B (FI) 95.89 65.35 43.40 0.00 0.00 97.27 0.00 43.13 1684.16
1 demo
22 Gemma-2B (EstTrue-Acc) 35.57 34.45 34.84 31.76 0.00 26.35 34.34 140.95
23 Gemma-2B (Systematic) 32.03 34.45 25.25 26.81 0.00 6.07 27.49 84.32
24 Gemma-2B (FI) 90.03 99.98 72.47 84.41 0.00 23.03 80.05 74.99 611.87
5 demos
25 Gemma-2B (EstTrue-Acc) 38.56 38.90 38.49 36.72 36.52 38.97 37.25 0.94
26 Gemma-2B (Systematic) 38.56 38.90 38.49 36.72 36.08 38.29 37.20 0.99
27 Gemma-2B (FI) 100.00 100.00 100.00 100.00 98.78 98.26 99.86 99.57 0.45
Table 18: Supplementing demonstrations, repeating format instructions, and extra fine-tuning with formats’ data reduce format bias. Performance of ChatGPT and Gemma-2B-it (without CoT) on MMLU. All results are reliably measured by Theorem 3.1.
In this section, we present the numerical results of our proposed techniques for mitigating format biases using ChatGPT on MMLU, as shown in Tab. 18 indexes 1-15.
  1. 1.
    Demonstrations with formats reduce bias (Indexes 7-12). From Tab. 18 indexes 7-12, we observe that using demonstrations with formats generally increases the average of FI scores, from $72.99\%$ without any demonstration (index 3), to $79.20\%$ with using one demonstration and $85.32\%$ with using $5$ demonstrations. Moreover, we find that the performance does not scale linearly with the FI score, indicating that simply increasing the FI score does not necessarily improve the models’ performance or reduce format biases.
  2. 2.
    Repeating format instructions reduces format bias (Indexes 4-6). From Tab. 18 index 6, most of the formats, repeating the format instruction can increase the FI score (compared to index 3), except for the “Placeholder”. Manual investigation reveals that repeatedly using the “Placeholder” format confuses the model about the actual location of the placeholder, leading to the model omitting the format. Nevertheless, this strategy generally reduces the format bias by decreasing the variance of results from formats other than “Placeholder”, leading to overall reduction.
  3. 3.
    Fine-tuning with additional format data can eliminate format bias (Indexes 13-15). Finetuning mostly eliminates the format bias problem of the LLM with the bias score only $0.71\%^{2}$ from Tab. 18 indexes 13-15, while increasing the average FI score up to almost perfect with $99.59\%$. This demonstrates that finetuning can help LLMs become more familiar with format tokens and requirements, reducing bias towards different formats.
For Gemma-2B-it model indexed 16-27, we observe that:
  1. 1.
    Repeating format instruction (Indexes 19-21). This approach does not resolve Gemma-2B-it’s inability to adhere to the “Brackets” and “Parentheses” instructions. However, disregarding these two formats, the strategy effectively reduces format bias in the model’s performance.
  2. 2.
    Demonstrations (Indexes 22-27). Demonstrations significantly mitigate format bias, as evidenced by higher FI scores and reduced variance across format performances.
Appendix D Experimental Details
D.1 Dataset Details
We provide descriptions of all datasets we use in this paper.
MMLU Hendrycks et al. (2021).
MMLU is a benchmark for evaluating the performance of language models on Multiple Choices Question on a wide range of subjects across STEM, the humanities, social sciences, and other areas, testing the model’s ability to understand and reason in diverse domains.
BBH Suzgun et al. (2023).
BBH is a MCQ dataset which includes a variety of challenging benchmarks that require advanced reasoning, comprehension, and other complex cognitive skills.
GSM8K Cobbe et al. (2021).
GSM8K is a dataset of 8,000 math word problems designed for grade school students. The problems require not just basic arithmetic but also multi-step reasoning to solve.
HotpotQA Yang et al. (2018).
HotpotQA is a question-answering dataset with a focus on multi-hop reasoning. It contains questions that require finding and combining information from multiple Wikipedia articles to derive the answer.
FairytaleQA Xu et al. (2022).
FairytaleQA is a dataset designed for evaluating narrative comprehension, particularly in the context of children’s fairytales. It includes questions that test understanding of characters, plots, and settings in fairytales.
SciDocsRR Cohan et al. (2020).
SciDocsRR is a dataset for evaluating information retrieval systems, particularly in the scientific domain. It includes tasks like citation prediction, document classification, and other retrieval-based evaluations.
SemEval2017 Augenstein et al. (2017).
SemEval2017 is part of an ongoing series of evaluations for semantic analysis in natural language processing. It includes a wide range of tasks such as sentiment analysis, semantic textual similarity, and information extraction.
SciREX Jain et al. (2020).
SciREX is a dataset for evaluating models on the task of information extraction from scientific literature. It focuses on extracting entities, relations, and other structured information from research papers.
D.2 Experimental Results
We present the hyperparameters setting for our experiments below.
Gemma-7B-it Team et al. (2024).
For Gemma 7B-it, use the weights from Google and Huggingface [^2] https://huggingface.co/google/gemma-7b-it. We use Nucleus Sampling Holtzman et al. (2020) as our decoding strategy with a $p$ value of $0.95$, a temperature value of $0.1$, and a window size of $1024$.
Mistral-7B-it-v0.2 Jiang et al. (2023).
For Mistral 7B-it, use the weights from MistralAI and Huggingface [^3] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2. We use Nucleus Sampling Holtzman et al. (2020) as our decoding strategy with a $p$ value of $0.9$, and a window size of $1024$.
ChatGPT (gpt3.5-turbo-0125) OpenAI (2022).
For ChatGPT, we use the system role: “You are helpful assistant!”. We set the “max_tokens” to be 1024, “top_p=1”, “frequency_penalty=0”, “presence_penalty=0”, and the model mode is “gpt3.5-turbo-0125”.
Datasets for finetuning ChatGPT and finetuning setups.
We preprocess the “auxiliary_train” [^4] https://huggingface.co/datasets/cais/mmlu/viewer/auxiliary_train dataset of MMLU Hendrycks et al. (2021), resulting in the training set of $6500$ samples as discussed in Section 6. We preprocess a small, distinct validation set with the same ratio as the training set among formats ”20-20-40-40-50-20-50”, resulting in a total of $240$ samples for validation.
We use the default finetuning setup of OpenAI for ChatGPT. Our finetuning costs $63.86$ US\$.
Appendix E Prompting
E.1 MCQ Prompt Details
The input for the models is the combination of the following components:
[…]
where non-CoT/CoT Instruction shows that model uses Zero-shot or Chain-of-Thought, given that
  • non-CoT $\times$ Char. Instruction = “Answer the following multiple-choice question by outputting only the designated character identifier.”
  • non-CoT $\times$ Text. Instruction = “Answer the following multiple-choice question by outputting the textual value of your choice without the character identifier without any textual description.”
  • CoT $\times$ Char. Instruction = “Answer the following multiple-choice question step-by-step by outputting only the designated character identifier.”
  • CoT $\times$ Text. Instruction = “Answer the following multiple-choice question step-by-step by outputting the textual value of your choice without the character identifier.”
Question is the main content of the task and CoT Wrapping is wrapping instruction if using CoT. i.e. CoT Wrapping = “Wrap your final answer by <ANSWER> and </ANSWER>.”
E.2 Wrapping Prompt Details
The input for the models is the combination of the following components:
[…]
where non-CoT/CoT Instruction shows that model uses Zero-shot or Chain-of-Thought, given that
  • If MCQ task (MMLU,BBH)
    1. 1.
      non-CoT Instruction = “Answer the following multiple-choice question by outputting only the designated character identifier.”
    2. 2.
      CoT Instruction = “Answer the following multiple-choice question step-by-step by outputting only the designated character identifier.”
  • If generation task (GSM8K, HotpotQA, FairytaleQA)
    1. 1.
      non-CoT Instruction = “Answer the following question.”
    2. 2.
      CoT Instruction = “Answer the following question step by step.”
Question is the main content of the task, and Wrapping Format Instruction is the format we want the model to output, detailed as
  • Special Character wrapping = “Wrap your final answer by <ANSWER> and </ANSWER>.”
  • Bolding wrapping = “Wrap your final answer in bold by enclosing it with double asterisks.”
  • Italicizing wrapping = “Wrap your final answer in italics by enclosing it with single asterisks.”
  • Brackets wrapping = “Wrap your final answer using double square brackets.”
  • Parentheses wrapping = “Wrap your final answer using double parentheses.”
  • Placeholder wrapping = “Wrap your final answer by filling in the placeholder below: ‘So the answer is: {{placeholder}}”’
  • Quoting wrapping = “Wrap your final answer using triple quotation marks.”
E.3 List Prompt Details
For dataset SciDocsRR, the input for the models is the combination of the following components:
[…]
where
  • Information = “Given a query, and a list of documents:
     Topic: Topic.
    List of documents: Samples
  • Requirement = “You are required to transform the list of documents into a binary list of $1$ or 00 where $1$ indicates the document brings useful information to the topic, and 00 indicates the document does not bring useful information to the topic.”
  • List Format Instruction includes four categories:
    1. 1.
      Python = “Generate your binary list as a Python list”
    2. 2.
      Bullet = “Generate your binary list using bullet points”
    3. 3.
      Special Character = “Generate your binary list using <SEP> to separate elements”
    4. 4.
      New Line = “Generate your binary list such that each element is in a new line”
  • non-CoT / CoT Instruction includes:
    1. 1.
      non-CoT Instruction = “without any explanation.”
    2. 2.
      CoT Instruction = “step by step”
  • CoT Wrapping = “Wrap your final list by <ANSWER> and </ANSWER>.”
For dataset SemEval2017, the input for the models is the combination of the following components:
[…]
where
  • Requirement = “Extract a list of keyphrases from the following document:”
  • Document is the main content of the task.
  • List Format Instruction includes four categories:
    1. 1.
      Python = “Generate your binary list as a Python list”
    2. 2.
      Bullet = “Generate your binary list using bullet points”
    3. 3.
      Special Character = “Generate your binary list using <SEP> to separate elements”
    4. 4.
      New Line = “Generate your binary list such that each element is in a new line”
  • non-CoT / CoT Instruction includes:
    1. 1.
      non-CoT Instruction = “without any explanation.”
    2. 2.
      CoT Instruction = “step by step”
  • CoT Wrapping = “Wrap your final list by <ANSWER> and </ANSWER>.”
E.4 Mapping Prompt Details
For all three datasets, we use the following formula for the input of the models
[…]
where
  • Requirement = “Extract the entities reflecting the tasks in the following document:” if using non-CoT model and “Extract the entities reflecting the tasks in the following document step-by-step:” if using CoT model
  • Document is the main content of the task.
  • CoT Wrapping = “Wrap your final list by <ANSWER> and </ANSWER>.”
  • Mapping Format Instruction starts with defining a specific format for the model and then instructs the model to follow. In detail, we have
    1. 1.
      For Easy dataset, we define:
      1 JSON_FORMAT = {
      2 ’’Task’’: [...]
      3 }
      1 YAML_FORMAT = ’’’’’’ Task: [...] ’’’’’’
      Then
      • JSON Mapping = “Your output must be a Python dictionary with the key ‘Task’ and value as a list of task name entities: {str(JSON_FORMAT)}
      • YAML Mapping = “Your output must be in YAML format: {str(YAML_FORMAT)}
    2. 2.
      For Medium dataset, we define:
      1 JSON_FORMAT = {
      2 ’’Task’’: [...],
      3 ’’Method’’: [...]
      4 }
      1 YAML_FORMAT = ’’’’’’
      2 Task: [...]
      3 Method: [...]
      4 ’’’’’’
      Then
      • JSON Mapping = “Your output must be a Python dictionary with the keys ‘Task’ and ‘Method’, and value is a list of task name entities and method name entities: {str(JSON_FORMAT)}
      • YAML Mapping = “Your output must be in YAML format: {str(YAML_FORMAT)}
    3. 3.
      For Hard dataset, we define:
      1 JSON_FORMAT = {
      2 ’’Task’’: [...],
      3 ’’Method’’: [...],
      4 ’’Material’’: [...],
      5 ’’Metric’’: [...]
      6 }
      1 YAML_FORMAT = ’’’’’’
      2 Task: [...]
      3 Method: [...]
      4 Material: [...]
      5 Metric: [...]
      6 ’’’’’’
      Then
      • JSON Mapping = “Your output must be a Python dictionary with the keys are ‘Task’, ‘Method’, ‘Material’, ‘Metric’, and value is a list of task name entities, method name entities, material name entities, metric name entities: {str(JSON_FORMAT)}
      • YAML Mapping = “Your output must be in YAML format: {str(YAML_FORMAT)}